OpenStudy (anonymous):

Factor the trinomial and write each factor as a polynomial in descending order: 3x^2-2x-21

OpenStudy (aroub):

(3x-7)(x+3)

myininaya (myininaya):

i would look for factors of 3(-21) that add up to be -2 3(-21)=9(-7)=-9(7) and -9+7=-2 so we have 3x^2-9x+7x-21 3x(x-3)+7(x-3) (x-3)(3x+7)

OpenStudy (aroub):

sorry ,( 3x+7)(x-3)

OpenStudy (anonymous):

(3x-7)(x+3)

myininaya (myininaya):

it is (x-3)(3x+7) or (3x+7)(x-3)

OpenStudy (anonymous):

x1=3 or x2=-7/3 ax^2+bx+c= a.(x-x1).(x-x2) Then 3x^2-2x-21= 3(x-3)(x+7/3)=(x-3)(3x+7)